Solve the initial value problem. y – 6y +9y = 28(t – 1/3) cost y(0) = 0, y(0) = 0
Answer #1

constants. Į so, Sol? y - Gy +94 = 28(+-) cost ; 7(0) = 0 = y(0) Taking Laplace transform, L18” - 48 + 3) = L 2 ( + - A) a8-21 1+1 !!= 1 2 = 1+1 Now, Let F(x) = -,2 ; 1{F(x)} = f(t); 512=a} = {{c} {E} {apply chain yang = the at fox ma41,2,3...)